OpenStudy (anonymous):

evaluate x/y for x=2/3 and y=6/7

OpenStudy (anonymous):

do you mean (2/3)/(6/7) ? if so than take the recipicle of the second number flip it upside down and multiply out 2/3*7/6= 14/18

OpenStudy (anonymous):

no i do not

OpenStudy (amoodarya):

2 -- 3 2*7 ----= ----- 6 3*6 -- 7 far by far , near by near